### v1

# Relating Checkpoint Update Probabilities to Momentum Parameters in Single-Loop Variance Reduction Methods

Abstract
Variance reduction (VR) is a crucial tool for solving finite-sum optimization problems, including the composite general convex setting, which is the focus of this work. On the one hand, denoting the number of component functions as $n$ and the target accuracy as $\epsilon$, some VR methods achieve the near-optimal complexity $\widetilde{\mathcal{O}}\left(n+\sqrt{n}/\sqrt{\epsilon}\right)$, but they all have complicated nested structure and fail to provide convergence guarantee for the iterate sequence itself. On the other hand, single-loop VR methods, being free from the aforementioned disadvantages, have complexity no better than $\mathcal{O}\left(n+n/\sqrt{\epsilon}\right)$ which is the complexity of the deterministic method FISTA, thus leaving a critical gap unaddressed. In this work, we propose the Harmonia technique which relates full-gradient point (i.e., checkpoint) update probabilities to momentum parameters in single-loop VR methods. Based on this technique, we further propose to vary the growth rate of the momentum parameter, creating a novel continuous trade-off between acceleration and variance reduction, controlled by the key parameter $\alpha\in[0,1]$. The proposed techniques lead to following favourable consequences. First, several known complexity of quite different algorithms are re-discovered under the proposed unifying algorithmic framework Katyusha-H. Second, under an extra mild condition, Katyusha-H achieves the near-optimal complexity for $\alpha$ belonging to a certain interval, highlighting the effectiveness of the acceleration-variance reduction trade-off. Moreover, by choosing smaller $\alpha$, Katyusha-H can achieve the same complexity (in the big-$\mathcal{O}$ sense) with less per-iteration cost. Last, without extra conditions, Katyusha-H achieves the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ with $\alpha=1$ and proper mini-batch sizes. The proposed idea and techniques may be of general interest beyond the considered problem in this work.
keywords:
finite-sum optimization, single-loop methods, variance reduction, acceleration, Katyusha momentum
1 Introduction
In this work, we consider the following composite general convex optimization problem:
$\min_{x\in\mathbb{R}^{d}}F(x):=f(x)+l(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)+l(x),$ (1)
where $f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)$ with $f_{i}(x)$ being convex and smooth, and $l(x)$ is convex, lower semicontinuous (but possibly non-differentiable) function which admits an efficient proximal operator. Problems of this type are ubiquitous in statistics and machine learning such as regularized empirical risk minimization [1].
When $n$ is large, the computation of $\nabla f(x)$ is expensive since it requires computing $n$ component gradients. Gradient-based methods (e.g., the forward-backward splitting algorithm [2] and its inertial version [3, 4]) require access to the full gradient at every iteration, hence are unpractical. By utilizing the finite-sum structure of (1), stochastic gradient methods use the gradient of one or a few component(s) $\nabla f_{i}$ to construct the stochastic gradient estimator. For example, stochastic gradient descent (SGD) uses $\nabla f_{i}(x)$ (ii is chosen from $n]:=\{1,2,\cdots,n\}$ according to the discrete uniform distribution) as an unbiased estimator of $\nabla f(x)$ in each iteration, hence enjoys a low per-iteration cost. However, due to the non-diminishing variance of its gradient estimator, the convergence speed of SGD is very slow. Even when the objective is smooth and strongly convex, SGD can only achieve sublinear convergence.
Variance reduction (VR) methods [5, 6, 7, 8, 9, 10, 11, 12] utilize history information to construct variance-vanishing stochastic gradient estimators, enjoying both low per-iteration cost and fast convergence speed. In fact, the convergence rate of VR methods improves significantly upon SGD, and matches that of traditional gradient-based methods. For example, VR methods achieve the linear convergence rate in the smooth strongly convex problems.
[…]
where $i_{t}$ is chosen from $[n]$ according to the discrete uniform distribution at $t$th iteration. Then the iterate is iterated with the formula $x_{t+1}=x_{t}-\eta\tilde{\nabla}f_{i_{t}}(x_{t})$, where $\eta>0$ is the step size. After $m$ iterations, the checkpoint is updated (e.g., it can be chosen as the average of last $m$ iterates) and the aforementioned procedure is repeated. Hence the vanilla SVRG method, along with many other VR methods [13, 14, 7, 15], have nested algorithmic structure.
VR methods achieve well-known success in solving the composite general convex problem (1) under the incremental first-order oracle (IFO) framework [16], which evaluates the efficiency of an algorithm by the number of evaluating component gradients $\nabla f_{i}(x)$ needed to achieve an $\epsilon$-accurate solution. The vanilla SVRG relies on the strong convexity of $f$ to combine different outer iterations in the proof, hence cannot directly solve general convex problem theoretically. SVRG++ [17] resolves this issue by introducing the technique of doubling the inner loop length (i.e., $m_{s}=2^{s}\cdot m_{0}$ where $m_{s}$ is the inner loop length of the $s$th outer iteration), achieving the IFO complexity $\mathcal{O}(n\log(1/\epsilon)+1/\epsilon)$.
Katyushans [8] uses the Katyusha momentum [8] to mitigate possible negative effects of Nesterov acceleration in stochastic setting. It achieves the complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$, and the Katyusha momentum turned out to be instrumental for the acceleration of VR methods in the composite general convex problem. MiGnsc[18] simplifies the iteration structure of Katyushans and achieves essentially the same complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$. Using the Katyusha momentum in inner iterations, ASVRG [19] achieves the complexity $\mathcal{O}(n\log(1/\epsilon))+\sqrt{n}/\sqrt{\epsilon})$, which nearly matches the theoretical lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ [20]. By utilization of the Katyusha momentum, the acceleration techniques in [21, 22], and a multi-stage inner loop length selection scheme, Varag [23] also achieves the complexity $\mathcal{O}(n\log(n)+\sqrt{n}/\sqrt{\epsilon})$. VRADA [24] incorporates the Katyusha momentum with a novel initialization technique, achieving the complexity $\mathcal{O}(n\log\log(n)+\sqrt{n}/\sqrt{\epsilon})$. Based on these methods, some researchers employ the AdaGrad idea [25] to compute step sizes [26], maintaining similar complexity under additional assumptions such as the compactness of the underlying set which the optimization variable belongs to.
As introduced above, some VR methods can achieve the near-optimal complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$, where the $\widetilde{\mathcal{O}}$ notation hides logarithmic factors. However, all such methods have the nested algorithmic structure, which brings two major drawbacks. First, this complicated structure renders the implementation of the algorithm more troublesome. Second, these methods only has convergence guarantee for the checkpoint sequence. The checkpoint is usually not the last iterate of the last $m$ inner iterations [^1] Theoretically, choosing the last iterate of last $m$ inner iterations to be the new checkpoint deteriorates the complexity. Hence existing methods usually choose the checkpoint with certain (weighted) average, although this is empirically undesirable. and a new checkpoint can be obtained only after an entire outer iteration. For similar reasons, indirect methods [27], which adds strongly convex regularization terms to (1) with decreasing regularization parameters and uses VR methods to solve resultant strongly convex problems, are also undesirable.
In contrast, single-loop VR methods are easy to implement and typically provide convergence guarantee for the iterate sequence itself. However, the literature on single-loop VR methods for the composite general convex problem (1) is relatively limited. The complexity of single-loop VR methods are no better than $\mathcal{O}(n+n/\sqrt{\epsilon})$, which is exactly the complexity of FISTA [4]. Note that FISTA is an extension of Nesterov’s accelerated gradient method [28] for the composite optimization problem, and it requires full gradient computation in every iteration, hence is a gradient-based method.
In this line of works, L-SVRG [9] and L-Katyusha [9] are first single-loop VR methods. In L-SVRG, the checkpoint, denoted as $w_{t}$ (where $t$ is the iteration number), is updated in a randomized way. Concretely, at the end of the $t$th iteration, $w_{t+1}$ is replaced by the new iterate [^2] Once the replacement occurs, a full gradient is computed at $w_{t+1}$. with probability $p$ or remains $w_{t}$ otherwise, where $p\in[0,1]$ is a constant which takes a small value, e.g., $1/n$. However, L-SVRG and L-Katyusha require the smooth part $f$ to be strongly convex and the non-smooth part $l\equiv 0$, hence cannot solve the composite general convex problem (1). L2S [29] is a single-loop VR method that uses the SARAH gradient estimator [7] and a constant checkpoint update probability, having the complexity $\mathcal{O}(n+\sqrt{n}/\epsilon)$. However, it requires that $l\equiv 0$ in (1) and the convergence guarantee is only for the expected squared gradient norm, not the expected functional gap $\mathbb{E}[F(\cdot)-\inf F]$ which is usually adopted.
Without using the Katyusha momentum, Driggs et al. [30] proposed a universal single-loop acceleration framework for several frequently used variance reduced gradient estimators based on the Nesterov acceleration scheme. Under the composite general convex setting, their method has the convergence rate $\mathcal{O}(c_{n}/T^{2})$, where $T$ is the number of total iterations and $c_{n}$ is an $n$-dependent constant. Curiously, using mini-batch cannot improve the complexity of this scheme, regardless of which estimator is employed. For example, when the SVRG estimator is used, as the mini-batch size $b$ grows, $c_{n}$ is decreased but the per-iteration cost is increased. As a result, the gain in the iteration complexity and the loss in the per-iteration cost cancels perfectly, leading to the complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$ for any $b\in[n]$. The checkpoint update probability in [30] is also a constant $p=b/n$.
Note that setting $p$ to a constant is essential for the proof of [30], since Driggs et al. estimate the summation of the bias term and mean squared error of the gradient estimator respectively along the entire optimization process. The constant probability setting renders certain coefficient (in the proof) essentially the truncation of a geometric series hence upper-bounded and independent of the total iteration number $T$. If a non-constant probability setting is to be directly applied, a positive lower bound of the probabilities is likely to be required to prevent aforementioned coefficients from exploding. But in this way, the complexity arguably relates to the aforementioned lower bound in essential, hence cannot be improved in terms of the $n$-dependence.
After the completion of this manuscript, we came to realize the article [31] in which the SIFAR algorithm is proposed [^3] During the literature research, we thoroughly searched for all related articles of [24], which proposes the VRADA method enjoying the currently best complexity among nested VR methods, and [30], which is the only known single-loop VR method with momentum technique solving (1) to the best of our knowledge. The reason that we did not realize [31] is partly due to that [31] accidentally missed these articles.. The SIFAR algorithm is a single-loop VR method with acceleration technique and the checkpoint update probability $p_{t}$ varies between different stages. [31] denotes the iteration in which the checkpoint is updated for the first time as $t_{1}$, which follows the geometric distribution. Then there are three stages of SIFAR: $p_{t}\equiv 1/(n+1)$ for the first stage $0\leq t\leq t_{1}$, $p_{t}=4/(t-t_{1}+3\sqrt{n})$ for the second stage $t_{1}<t\leq t_{1}+n+3-3\sqrt{n}$, and $p_{t}\equiv 4/(n+3)$ for the third stage $t>t_{1}+n+3-3\sqrt{n}$.
The complexity $\mathcal{O}(n\min\{1+\log(1/(\epsilon\sqrt{n})),\log\sqrt{n}\}+\sqrt{(nL)/\epsilon})$ is claimed in [31]. However, we believe that the proofs in [31] are incorrect after careful consideration. First, Theorem $1$ of [31] is incorrect. In Lemma $3$ of [31], an upper bound is given for the quantity $\mathbb{E}\left\{[\eta_{t-1}/(p_{t-1}\theta_{t-1}^{2})][f(w_{t})-f(x^{*})]\right\}$ for any $t>t_{1}$. Since $t_{1}$ is a random variable (r.v.) that may take any positive integer value in the probability space, $t$ should also be a r.v. to satisfy the inequality $t>t_{1}$, hence the coefficient $\eta_{t-1}/(p_{t-1}\theta_{t-1}^{2})$ is a r.v. since it depends on $t$ and $t_{1}$. In [31], this coefficient is directly taken out of the expectation $\mathbb{E}$ and both sides of the inequality are multiplied by $(p_{t-1}\theta_{t-1}^{2})/\eta_{t-1}$ to obtain an upper bound on the concerned quantity $\mathbb{E}(f(w_{t})-f(x^{*}))$. However, this reasoning is illegal and the upper bounding of $\mathbb{E}(f(w_{t})-f(x^{*}))$ is elusive. Hence, Theorem $1$ of [31] is incorrect and the difficulty there is critical.
Consider two correlated r.v.’s $X_{1}$ and $X_{2}$, the estimating of terms of the form $\mathbb{E}(X_{1}X_{2})$ [^4] Sometimes such terms take the form of the expectation of a sum of products of correlated r.v.’s, i.e., $\mathbb{E}\sum_{t=0}^{T}X_{t}^{(1)}X_{t}^{(2)}$, where $X_{t}^{(1)}$ and $X_{t}^{(2)}$ are correlated r.v.’s for $0\leq t\leq T$. But we take the simple $T=0$ case for the ease of discussion. is of fundamental importance in the analysis of randomized methods. In some cases, we know how to estimate $\mathbb{E}(X_{2})$ or $\mathbb{E}(X_{2})$ is the concerned quantity, and we want to apply the known inequality or obtain $\mathbb{E}(X_{2})$ alone. For example, such terms arise when the step size is computed via the current and/or some previous (incremental) gradients, e.g., the step size is determined by the Barzilai-Borwein-type formulae [32] or the AdaGrad-type formulae [25]. Then whether one can successfully estimate such terms determines the success or failure of the convergence analysis [33]. And when estimating them are feasible, the way of doing the estimates critically affects the ultimate theoretical complexity [34]. Hence, the problem that Theorem $1$ of [31] faced is beyond a trivial mis-reasoning.
Second, Corollary $1$ of [31] is incorrect. Due to the fact that $t_{1}$ is a r.v., the iterations needed by the SIFAR algorithm to achieve an $\epsilon$-accurate solution is also a r.v. To the best of our knowledge, this is in contrast to other existing single-loop VR methods, where the required iterations is deterministic given the target accuracy $\epsilon$. Hence the analysis of the complexity should be conducted differently. For a single-loop VR method with probabilistic update of the checkpoint, denote the computational cost of the $t$th iteration as $C_{t}$ (note that $C_{t}$ is a r.v.), and denote the iterations needed to reach an $\epsilon$-accurate solution as $T(\epsilon)$. Then the total cost in expectation is $\mathbb{E}\sum_{t=0}^{T(\epsilon)}C_{t}$. When $T(\epsilon)$ is deterministic, we have $\mathbb{E}\sum_{t=0}^{T(\epsilon)}C_{t}=\sum_{t=0}^{T(\epsilon)}\mathbb{E}(C_{t})$. However, when $T(\epsilon)$ is a r.v., certain propositions, e.g., the Wald’s theorem or its generalized versions [35, 36], should be applied in such case (if they are applicable at all under the design and characteristics of the SIFAR algorithm).
In all, despite the appealing novelty and innovation of [31], we believe that the convergence analysis and complexity analysis concerning the general convex setting are incorrect, hence the corresponding complexity is invalid.
Inspired by the aforementioned theoretical and practical merit of the single-loop methods, we take this line of research further in this work. We propose the Harmonia technique which relates checkpoint update probability $p_{t}$ to momentum parameter $\alpha_{t}$ in single-loop VR methods to construct provably better algorithms. Our main contributions are summarized as follows:
  • Along with the existing Nesterov acceleration framework and the Katyusha momentum, we propose the Harmonia technique to accelerate single-loop VR methods in the composite general convex problem. With mini-batch size properly set, the proposed Katyusha-H method achieves the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$, which significantly improves upon the previous best $\mathcal{O}(n+n/\sqrt{\epsilon})$ of single-loop VR methods and matches the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ up to logarithmic factors.
  • Based on the Harmonia technique, we further propose to vary the growth rate of the momentum parameter, creating a novel continuous trade-off between acceleration and variance reduction, controlled by the key parameter $\alpha\in[0,1]$. Under an extra mild condition, Katyusha-H achieves $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ for $\alpha$ belonging to an explicitly given interval. In many cases, $\alpha=1$ which corresponds to the largest allowable acceleration strength does not belong to the aforementioned interval, hence showing the effectiveness of the proposed acceleration-variance reduction trade-off. Moreover, by choosing smaller $\alpha$, Katyusha-H maintains the near-optimal complexity while having less per-iteration cost.
  • Several known complexity of existing algorithms that use different optimization techniques and exhibit different structure are re-discovered under the unifying algorithmic framework Katyusha-H.
  • Based on the conducted theoretical analysis, we provide intuitive interpretations for the proposed $p_{t}-\alpha_{t}$ correspondence and the key parameter $\alpha$, which may be of general interest.
The rest of this article is organized as follows: Section 2 introduces notations, definitions and assumptions. Section 3 presents the proposed Katyusha-H method and related preliminary explanations. Section 4 presents our theoretical findings and detailed discussions and interpretations concerning the theoretical results, where most of the proofs are deferred to the Appendix. Section 5 summarises the proposed idea and technique from a high-level perspective and relist the contributions briefly.
2 Notations and Assumptions
For a differentiable function $f$, its gradient at $x$ is denoted by $\nabla f(x)$. The Euclidean norm of a vector $x\in\mathbb{R}^{d}$ is denoted as $\|x\|$. The set of all non-negative integers is denoted as $\mathbb{N}$, and the set of all positive integers is denoted as $\mathbb{N}^{+}$. The set $\{1,2,\cdots,n\}$ is denoted by $[n]$. The symbol $\mathbb{E}$ denotes the expectation of a random element, and $\mathbb{E}_{X}$ denotes an expectation over the randomness of a random variable $X$ while conditioning on all other random variables.
We define computational costs by making use of the IFO framework of [16], where we assume that sampling an index $i$ and computing the pair $(f_{i}(x),\nabla f_{i}(x))$ incurs a unit of cost. In the composite general convex setting, $x$ is called an $\epsilon$-accurate solution iff $\mathbb{E}\left\{F(x)-F(x^{*})\right\}\leq\epsilon$, where $x^{*}$ is a solution to the problem (1). The IFO complexity is the minimum IFO calls needed to reach an $\epsilon$-accurate solution.
In this work, the following standard assumptions will be made throughout:
A1 The optimization problem (1) has at least a solution, and we denote one fixed solution as $x^{*}$.
A2 Functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are $L$-smooth for some $L>0$:
$f_{i}(y)\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2},\quad\forall x,y\in\mathbb{R}^{d}.$
A3 Functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are convex:
$f_{i}(y)\geq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle,\quad\forall x,y\in\mathbb{R}^{d}.$
A4 The function $l$ is convex, lower semi-continuous (but possibly non-differentiable) function which admits an efficient proximal operator.
3 The Katyusha-H Method
The proposed Katyusha-H method is formally described in Algorithm 1. The three sequences formulation, consisting of ${x_{t}},{y_{t}},{z_{t}}$, is ubiquitous in accelerated gradient methods [37] and can be interpreted by the linear coupling interpretation [38] as follows. Consider a simplified setting where $l\equiv 0$ so that $F\equiv f$ and the mini-batch size $b=n$ so that $\tilde{g}_{t+1}=\nabla f(x_{t+1})$. Then the $\{z_{t}\}$ sequence is viewed and analysed as a mirror descent sequence, where an error term proportional to $\|\nabla f(x_{t+1})\|^{2}$ is introduced in its analysis. The $\{y_{t}\}$ iteration can be equivalently formulated as $y_{t+1}=y_{t}-\tau_{t}\alpha_{t}\eta\nabla f(x_{t+1})$, which further reduces to $y_{t+1}=y_{t}-\eta\nabla f(x_{t+1})$ when $\tau_{t}\alpha_{t}=1$. Hence $\{y_{t}\}$ is viewed and analysed as a gradient descent sequence, where an improvement proportional to $\|\nabla f(x_{t+1})\|^{2}$ is made in each iteration. Finally, $x_{t+1}$ linearly couples $z_{t}$ and $y_{t}$ to achieve acceleration, in the spirit of leveraging the aforementioned duality of the mirror descent and the gradient descent.
Algorithm 1 uses the mini-batch version of the SVRG estimator $\tilde{g}_{t+1}$, where the full gradient $\nabla f(w_{t})$ is required. The iterate $w_{t}$ is thus called the full gradient point, also called the checkpoint. In algorithm 1, $w_{t}$ is also involved in the coupling sequence through $x_{t+1}=\tau_{t}z_{t}+\xi w_{t}+(1-\xi-\tau_{t})y_{t}$, and the $w_{t}$ in this formula is known as the Katyusha momentum. By coupling with $w_{t}$, $x_{t+1}$ can be attracted to $w_{t}$ so that the variance of the estimator $\tilde{g}_{t+1}$ can be further reduced, enabling the acceleration technique to function well in the stochastic setting.
Algorithm 1 Katyusha-H
0: initial point $w_{1}=x_{1}=y_{1}=z_{1}$Parameters: step size $\eta$, mini-batch size $b$, Katyusha momentum parameter $\xi$, momentum parameter sequences $\{\alpha_{t}\}_{t\in\mathbb{N}}$, $\{\tau_{t}\}_{t\in\mathbb{N}^{+}}$
for $t=1,2,\cdots$ do
  $x_{t+1}=\tau_{t}z_{t}+\xi w_{t}+(1-\xi-\tau_{t})y_{t}$
  Pick a subset $J_{t}\subset[n]$ according to the discrete uniform distribution on the set of all subsets of $[n]$ which has the cardinality $b$
  $\tilde{g}_{t+1}=\frac{1}{b}\left(\sum_{j\in J_{t}}\nabla f_{j}(x_{t+1})-\nabla f_{j}(w_{t})\right)+\nabla f(w_{t})$
  $z_{t+1}=\arg\min_{z\in\mathbb{R}^{d}}\left\{\frac{1}{2\alpha_{t}\eta}\|z-z_{t}\|^{2}+\langle\tilde{g}_{t+1},z\rangle+l(z)\right\}$
  $y_{t+1}=x_{t+1}+\tau_{t}(z_{t+1}-z_{t})$
  $w^{t+1}=\begin{cases}y_{t}&\text{with probability }p_{t}\\ w_{t}&\text{with probability }1-p_{t}\end{cases}$
  $p_{t}=\frac{\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t}+\xi\alpha_{t}^{2}}{\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}}$
end for
The checkpoint $w_{t}$ is updated in a probabilistic way, which was first proposed by [9]. Most single-loop VR methods, including the pioneering work [9], use constant checkpoint update probability. To the best of our knowledge, the SVRLS method [39] is the only method that uses varying update probability $p_{t}=1/(at+1)$ where $0<a<1$. However, SVRLS is a non-accelerated method that is analysed by an AdaGrad-type regret analysis under an extra compact feasible set condition, requires that the nonsmooth part $l\equiv 0$ in (1), and unfortunately fails to provide convergence guarantee for the iterate sequence itself [^5] SVRLS thus does not have the last-iterate convergence guarantee..
The key feature of Algorithm 1 is the utilization of the proposed Harmonia technique, which is instrumental to the theoretical improvement of this work, and also leads to some novel insights into single-loop VR methods. These will be discussed concretely in the next section, where we present theoretical properties of the Katyusha-H method.
4 Main Results and Related Discussions
The following lemma shows that Algorithm 1 is well-defined and is fundamental to the convergence analysis of Algorithm 1. For example, it proves that in Algorithm 1 the parameter $p_{t}\in[0,1]$ hence is indeed a probability for all $t\in\mathbb{N}^{+}$, and that the coefficient of certain component of the Lyapunov function in the convergence analysis is non-negative for all $t\in\mathbb{N}^{+}$.
Lemma 1.
For $t\in\mathbb{N}$, define
$\alpha_{t}=\begin{cases}6,&0\leq t\leq 16\\ a_{\alpha}t^{\alpha},&t\geq 17\end{cases}\quad\text{with}\quad a_{\alpha}=\begin{cases}6,&\alpha=0\\ 1+\frac{\sqrt{2}}{4},&\alpha\in\left(0,1/2\right]\\ \frac{1}{3},&\alpha\in\left(1/2,3/4\right]\\ \frac{1}{4}\left(\frac{17}{16}\right)^{\alpha-1},&\alpha\in(3/4,1]\end{cases}.$
In Algorithm 1, set $\xi=1/(bc)$ where $b\in[n]$ is the mini-batch size and
$c=\max\left\{2,b^{-1}\max\left\{6/5,(1-\alpha_{17}^{-1})^{-1}\right\}\right\}+1,$
and $\tilde{\alpha}_{0}=\xi\alpha_{1}^{2}$. Then in Algorithm 1, $p_{t}\in[0,1]$ for $t\in\mathbb{N}^{+}$, $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq 0$ for $t\in\mathbb{N}$, and $\tau_{t},\xi,(1-\xi-\tau_{t})\in(0,1)$ for $t\in\mathbb{N}^{+}$. Moreover, there exists a constant $C>0$ which does not depend on any problem-dependent constants (e.g., $n$) such that [^6] As is shown in the proof, $C$ can take a mild value, e.g., $C=5$.
$c\leq C,\ \text{for all}\ \alpha\in[0,1].$
With Lemma 1, the convergence of Algorithm 1 can be guaranteed by the following theorem.
Theorem 1.
With the $\{\alpha_{t}\}$ sequence set as in Lemma 1, set $\tau_{t}=1/\alpha_{t}$ for $t\in\mathbb{N}^{+}$, $\eta\leq 1/(cL+L)$ and fix any $b\in[n]$. For all $T\in\mathbb{N}^{+}$, after $T$ iterations, Algorithm 1 produces a point $w_{T}$ satisfying the following bound:
$\alpha_{T}^{2}\mathbb{E}\left[F(y_{T+1})-F(x^{*})\right]+\left(\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{T}^{2}+\sum_{j=1}^{T}\alpha_{j}\right)\mathbb{E}\left[F(w_{T+1})-F(x^{*})\right]$
$+\frac{1}{2\eta}\mathbb{E}\|z_{T+1}-x^{*}\|^{2}\leq\alpha_{0}^{2}\left[F(y_{1})-F(x^{*})\right]+\tilde{\alpha}_{0}\left[F(w_{1})-F(x^{*})\right]+\frac{1}{2\eta}\|z_{1}-x^{*}\|^{2}.$ (2)
The setting of parameters of Algorithm 1 are determined in Lemma 1 and Theorem 1, and it is fairly intuitive. First, $\xi=1/(bc)$ is the coefficient of the Katyusha momentum, which is constructed to cancel certain terms in the upper bound of the variance of the SVRG estimator, i.e., to further reduce the variance (in the sense that one can use a tighter upper bound on the variance of the SVRG estimator with the presence of the Katyusha momentum). Hence it is expected that $\xi$ becomes smaller when the variance of the SVRG estimator decreases, which is actually the case since $\xi$ is non-increasing (and likely to decrease) when the mini-batch size $b$ increases. Second, the largest allowable step size $\eta=1/(cL+L)$ is non-decreasing (and likely to increase) when $b$ increases, which is reasonable since larger step size is expected to work when the variance of the stochastic estimator decreases.
Note that the convergence guarantee is for the iterate sequence $\{w_{t}\}$ itself [^7] By which we mean $F(w_{t})$ converges to $F(x^{*})$ in expectation. Convergence guarantee is also available for the $\{y_{t}\}$ sequence. but the convergence rate of $\{y_{t}\}$ is slower than that of $\{w_{t}\}$., hence the last-iterate convergence is available. In the right hand side (RHS) of (2), the coefficient $\alpha_{0}^{2}$ equals $36$ and is an absolute constant, $\tilde{\alpha}_{0}$ is upper bounded by $36$, and $1/(2\eta)$ equals $L(c+1)/2$ when taking the largest allowable step size, where $c\leq C$ by Lemma 1. Hence the convergence rate is $\mathcal{O}(1/(\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{T}^{2}+\sum_{j=1}^{T}\alpha_{j}))$ and is $n$-independent [^8] We refer to the proof of Theorem 2 in the appendix, where we show that $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{T}^{2}+\sum_{j=1}^{T}\alpha_{j}$ is lower bounded by a constant multiple of $T^{\alpha+1}$., which is in contrast to Theorem $5$ in [30], where the convergence rate has an $n$-dependent constant and the order of $n$ can be very large [^9] In Theorem $5$ of [30], using the same notation as [30], the dominant constant of the convergence rate is either $\mathcal{O}(\nu^{2})$ or $\mathcal{O}(c)$, which are both $n$-dependent..
It is interesting to notice the different roles of the checkpoint update probability $p_{t}$ and the mini-batch size $b$ in the convergence analysis. In expectation, one iteration of Algorithm $1$ costs $\mathcal{O}(np_{t}+b)$ IFO calls, hence $p_{t}$ and $b$ both relate to the per-iteration cost. In theory, $p_{t}$ and $b$ both relate to the variance of the SVRG estimator. When $p_{t}\equiv 1$ or $b=n$, the variance vanishes and the SVRG estimator reduces to the full gradient. However, in the convergence analysis, $p_{t}$ is chosen in accordance with $\alpha_{t}$ while $b$ can be chosen arbitrarily. This observation indicates that $p_{t}$ plays a more significant role in maintaining the convergence property.
With Theorem 1, one can obtain the number of iterations $T(\epsilon)$ needed to reach an $\epsilon$-accurate solution. We further estimate $\sum_{t=1}^{T(\epsilon)}\left(np_{t}+b\right)$ to obtain the IFO complexity of Algorithm 1.
Theorem 2.
Under the setting of Lemma 1 and Theorem 1, for any fixed $\alpha\in[0,1]$ in the definition of $\{\alpha_{t}\}_{t\in\mathbb{N}}$ in Lemma 1, the convergence rate of Algorithm 1 is $\mathcal{O}(1/T^{\alpha+1})$, the per-iteration IFO call in expectation is $\mathcal{O}(b+n/(bt^{1-\alpha})+n/t)$ for all $t\geq 17$. Without loss of generality, assume that the target accuracy $\epsilon\in(0,1)$, define $\hat{\alpha}=\log(2)/\log(\lceil 1/\epsilon\rceil)$, then the IFO call in expectation needed by Algorithm 1 to reach an $\epsilon$-accurate solution is
$\mathbb{E}C_{\text{comp}}=\begin{cases}\mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+n\log(1/\epsilon)\right),\quad&\alpha\in\left[0,\min\left\{\hat{\alpha},1/10\right\}\right]\\ \mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+\frac{n}{b}\frac{1}{\alpha}\frac{1}{\epsilon^{\alpha/(\alpha+1)}}+n\log(1/\epsilon)\right).\quad&\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right]\end{cases}$ (3)
In the statements above, $T\in\mathbb{N}^{+}$ is the total iteration number, $t\in[T]$ is the iteration counter, $n\in\mathbb{N}^{+}$ is the number of smooth component functions in (1), $b\in[n]$ is the mini-batch size.
We give some remarks on the threshold $\hat{\alpha}$. By the proof of Theorem 2, the complexity of the $\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right]$ case actually applies to all $\alpha\in\left(0,1\right]$. But to prevent the constant $1/\alpha$ from exploding when $\alpha\rightarrow 0^{+}$, more delicate proof is conducted where we introduce the threshold $\hat{\alpha}$ and discuss the two cases separately and carefully.
In the finite-sum optimization, one of the most fundamental topics is the trade-off between the iteration complexity and the per-iteration IFO cost, since essentially it is these two factors that determines the theoretical complexity of an optimization method. Handling this trade-off typically involves using different gradient estimators, using different optimization techniques, and even changing the algorithmic structure. For example, using full gradient gives the proximal gradient descent (PGD) a fast iteration convergence rate and high per-iteration IFO cost. While using the SGD estimator gives the proximal stochastic gradient descent (PSGD) contradictory characteristics. On the algorithmic structure side, note that the majority of variance reduction methods are nested, while PGD and PSGD are single-loop methods.
However, based on the proposed $p_{t}-\alpha_{t}$ correspondence, we can achieve this trade-off under the unifying framework of Algorithm 1, by changing the growth rates of the $\{\alpha_{t}\}_{t\in\mathbb{N}}$ sequence, which is novel to the literature to the best of our knowledge.
As proved in Theorem 2, less aggressive momentum acceleration (i.e., smaller $\alpha$) accompanies with smaller checkpoint update probability (i.e., smaller $p_{t}$), hence slower convergence rate can be compensated with lower per-iteration IFO cost in Algorithm 1, showing a novel continuous [^10] By continuous, we mean that the parameter $\alpha$ is allowed to take any value in the interval $[0,1]$. trade-off between the convergence rate and the per-iteration cost. The key parameter $\alpha\in[0,1]$ is thus a weight that controls the proposed acceleration-variance reduction trade-off, through which Algorithm 1 essentially controls the trade-off between the convergence rate and the per-iteration cost. Two consequences of this mechanism are listed as follows.
First, it is interesting to note that Algorithm 1 recovers some existing IFO complexity of very different algorithms under a unifying framework. Consider the $b=1$ case. When $\alpha=0$ (i.e., $\alpha_{t}\equiv 6$ for all $t\in\mathbb{N}$), the complexity of Algorithm 1 is $\mathcal{O}(n\log(1/\epsilon)+1/\epsilon)$, recovering the complexity of SVRG++ while removing some requirements concerning the knowledge of the solution $x^{*}$ needed by SVRG++. When $\alpha=1$ (i.e., $\alpha_{t}=\mathcal{O}(t)$), the complexity of Algorithm 1 is $\mathcal{O}(n+n/\sqrt{\epsilon})$, recovering the complexity of FISTA. Note that SVRG++ is a nested algorithm, uses the SVRG estimator, and employs the doubling technique, while FISTA is a single-loop algorithm, uses the full gradient as the iteration direction, and employs the Nesterov acceleration technique.
Second, since $\alpha$ in Lemma 1 belongs to $[0,1]$, we discover a series of novel complexity, from which one can choose a suitable $\alpha$ value to obtain near-optimal method according to the knowledge of $n$ and $\epsilon$ under a mild condition. Consider the $b=1$ case for Algorithm 1. When $n=10^{4},\epsilon=10^{-12}$, the complexity of SVRG++, FISTA, Algorithm 1 ($\alpha=1/2$), and the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ are $\mathcal{O}(10^{12})$, $\mathcal{O}(10^{10})$, $\mathcal{O}(10^{8})$ and $\mathcal{O}(10^{8})$, respectively. In fact, given $n\in\mathbb{N}^{+}$ and $\epsilon\in(0,1)$ satisfying
$n<1/\epsilon,$ (4)
for prefixed constants $C_{1}\in\left[1,+\infty\right)$ and $C_{2}\in\left[1,+\infty\right)$ (e.g., $C_{1}=C_{2}=2$), define [^11] We require that the choice of $C_{2}$ makes $\Delta_{2}$ strictly positive. Under the condition (4), this requirement can always be satisfied by choosing $C_{2}$ sufficiently close to $1$.
$\Delta=[\Delta_{1},\Delta_{2}]=\left[\frac{1-\frac{\log(n)}{\log(1/\epsilon)}-2\frac{\log(C_{1})}{\log(1/\epsilon)}}{1+\frac{\log(n)}{\log(1/\epsilon)}+2\frac{\log(C_{1})}{\log(1/\epsilon)}},\frac{1-\frac{\log(n)}{\log(1/\epsilon)}+2\frac{\log(C_{2})}{\log(1/\epsilon)}}{1+\frac{\log(n)}{\log(1/\epsilon)}-2\frac{\log(C_{2})}{\log(1/\epsilon)}}\right],$
one can always choose $\alpha\in[0,1]$ satisfying
$\alpha\in\begin{cases}\left[\max\left\{0,\Delta_{1}\right\},\Delta_{2}\right]&\quad\text{if}\ \Delta_{2}\leq\min\left\{\hat{\alpha},1/10\right\},\\ \left[\max\left\{\min\left\{\hat{\alpha},1/10\right\},\Delta_{1}\right\},\min\left\{1,\Delta_{2}\right\}\right]&\quad\text{if}\ \Delta_{2}>\min\left\{\hat{\alpha},1/10\right\},\end{cases}$ (5)
then direct calculation gives
$\frac{1}{\epsilon^{1/(\alpha+1)}}\leq C_{1}\frac{\sqrt{n}}{\sqrt{\epsilon}},\ \frac{n}{\epsilon^{\alpha/(\alpha+1)}}\leq C_{2}\frac{\sqrt{n}}{\sqrt{\epsilon}}.$
Hence the complexity in expectation of Algorithm 1 matches the theoretical lower bound up to a logarithmic factor under the $\alpha$ setting (5). This observation can be summarised in the following proposition.
Corollary 1.
Under the setting and assumptions of Theorem 2, assume that condition (4) holds, by choosing $b=1$ and $\alpha$ satisfying (5), the complexity in expectation of Algorithm 1 is $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$.
Proof.
Since $\Delta_{1}<\Delta_{2}$ and $\Delta_{2}>0$ under the condition (4) and the appropriate $C_{2}$ choice described in the footnote, the interval $\Delta$ is well-defined, and the intervals in (5) are well-defined and contained in $[0,1]$. Direct calculation shows that $\alpha\geq\Delta_{1}$ implies $1/\epsilon^{1/(\alpha+1)}\leq C_{1}\sqrt{n}/\sqrt{\epsilon}$ and $\alpha\leq\Delta_{2}$ implies $n/\epsilon^{\alpha/(\alpha+1)}\leq C_{2}\sqrt{n}/\sqrt{\epsilon}$. If $\Delta_{2}\leq\min\left\{\hat{\alpha},1/10\right\}$, under the first choice in (5), we have $\mathbb{E}C_{\text{comp}}$ takes the first form in (3) and $1/\epsilon^{1/(\alpha+1)}\leq\sqrt{n}/\sqrt{\epsilon}$, hence $\mathbb{E}C_{\text{comp}}=\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$. Similarly, one sees that if $\Delta_{2}>\min\left\{\hat{\alpha},1/10\right\}$, we have $\mathbb{E}C_{\text{comp}}=\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ under the second $\alpha$ choice in (5). In all, under the setting of this corollary, $\alpha$ satisfying (5) always exists and such choice of $\alpha$ yields the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$. ∎
For ease of discussion, we take $C_{2}=1$ in Corollary 1 and assume that $n$ in (1) is given. Some comments on Corollary 1 are listed as follows. First, we give intuitive explanations for the choice (5). The largest possible interval in (5) under the $\Delta_{2}\leq\min\left\{\hat{\alpha},1/10\right\}$ case is $\left[0,\min\left\{\hat{\alpha},1/10\right\}\right]$, and the largest possible interval in (5) under the other case is $\left[\min\left\{\hat{\alpha},1/10\right\},1\right]$, which is separated from the former interval by $\min\left\{\hat{\alpha},1/10\right\}$. Since $n$ is given and $C_{2}=1$, $\Delta_{2}$ is smaller when $\log(n)/\log(1/\epsilon)$ is larger, i.e., $\epsilon$ is larger. Hence the selection rule (5) essentially suggests that we should use mild acceleration and small update probability when the target accuracy is low (i.e., $\epsilon$ is relatively large), and we should use aggressive acceleration and large update probability when the target accuracy is high.
For ease of understanding, note that when the update probability $p_{t}$ is small, the per-iteration IFO cost is low and the variance of the gradient estimator $\tilde{g}_{t+1}$ is high, hence we refer to Algorithm 1 under small $p_{t}$ as the stochastic regime. Similarly, we refer to Algorithm 1 under large $p_{t}$ as the pseudo-deterministic regime [^12] Equivalently, Algorithm 1 with small $\alpha$ is the stochastic regime, and it is the pseudo-deterministic regime when $\alpha$ is large.. To understand the proper choice of regimes under different scales of the target accuracy, we turn to the following analogy.
Recalling the familiar algorithms PSGD and PGD, it is well known that PSGD is more suitable for solving (1) when the target accuracy is low (i.e., $\epsilon$ is large), since PSGD has low per-iteration cost and can make rapid progress in the early stage of optimization. However, when $\epsilon$ is small, PSGD has bad $\epsilon$ dependence in the complexity and converges very slowly, which is essentially due to the non-diminishing variance of the gradient estimator. Hence PGD can be favourable for solving (1) to a high target accuracy.
By analogy, for Algorithm 1, it is fairly reasonable to use the stochastic regime when $\epsilon$ is large, and use the pseudo-deterministic regime when $\epsilon$ is small. Moreover, it is well-known that the variance of the gradient estimator should be reduced to certain degree to make acceleration technique effective in randomized algorithms [^13] For example, [40, 41] use a fairly large mini-batch to make the Nesterov acceleration effective. For another example, the Katyusha momentum is proposed to cancel certain terms that come from a tight estimate of the variance of the SVRG estimator, so that the Nesterov acceleration technique can be utilized to accelerate SVRG in terms of the IFO complexity.. Hence, in the stochastic regime, mild acceleration can be added since the variance of the gradient estimator is mildly reduced, while aggressive acceleration can be applied in the pseudo-deterministic regime since the variance is significantly reduced.
Quantitatively, the important issue is exactly how much acceleration (controlled by $\alpha_{t}$) can the method adopt, given certain level of variance reduction (controlled by $p_{t}$). The answer to this question is provided by the proposed $p_{t}-\alpha_{t}$ correspondence, which essentially determines how much variance reduction should be conduct to perform certain level of acceleration, or vise versa.
For a clear interpretation, in the above analysis, the acceleration side and the variance reduction side is separated, and acceleration is regarded as an optional choice based on the existing level of variance reduction. While in Algorithm 1, the correspondence of the strength of acceleration and that of variance reduction comes simultaneously through the $p_{t}-\alpha_{t}$ correspondence, hence the name Harmonia [^14] Harmonia is the name of the Greek goddess of harmony and concord. for the proposed technique.
Second, Corollary 1 proves that more aggressive acceleration (i.e., larger $\alpha_{t}$) does not necessarily lead to better complexity in Algorithm 1 under the desirable constraint b=1b=1151515By desirable, we mean that smaller per-iteration cost is desirable if such choice does not deteriorate the complexity., while it is the trade-off between acceleration and variance reduction that actually matters. Since we take $C_{2}=1$, then $\Delta_{2}<1$ under the condition (4), which means that the largest allowable acceleration $\alpha_{t}=\mathcal{O}(t)$ (which corresponds to $\alpha=1$) can not provide best achievable complexity in either cases of (5) [^16] Note that $C_{2}=1$ is a desirable mild constant factor in the complexity. One can also easily check that $\Delta_{2}<1$ for a wide range of $\epsilon$ when $C_{2}>1$ takes a mild constant value..
This phenomenon somehow contradicts with the conventional choice in acceleration methods. Typically, since the convergence rate of an acceleration method is $1/\alpha_{T}^{2}$ where $T$ is the total iteration number, larger $\alpha_{t}$ setting is favourable as long as it is allowable. For example, in stochastic optimization problems where the gradient oracle contains stochastic noise, large momentum parameters may fail since the complexity explodes or the convergence fails. It is in such cases that the momentum parameter is then reduced in order to adapt to the inexactness of the gradient estimator[42].
In contrast, all $\alpha\in[0,1]$ settings are allowable in Algorithm 1, but we actively choose to reduce the acceleration strength and successfully get compensated by the reduction of the per-iteration cost in expectation. As a result, Corollary 1 proves that such trade-off can be utilized to get near-optimal complexity in expectation, while adopting the most aggressive acceleration under high-cost-based variance reduction is theoretically inferior.
Third, when the target accuracy $\epsilon$ can be determined in advance and the condition (4) is satisfied, the parameter setting (5) has its own merits. Concretely, since the mini-batch size $b=1$ is the smallest possible choice and $p_{t}$ is smaller than that under larger $\alpha$ choice (e.g., $\alpha=1$), Algorithm 1 under (5) achieves the near-optimal complexity with less per-iteration cost, which is desirable in the finite-sum optimization.
We believe that the proposed active acceleration-variance reduction trade-off is of theoretical value, and may be of general interest beyond the considered problem setting and methods of this article.
Finally, without requiring the condition (4) and the knowledge of $\epsilon$, Algorithm 1 achieves the near-optimal complexity in expectation, which significantly improves upon the previous best $\mathcal{O}(n+n/\sqrt{\epsilon})$ complexity of single-loop VR methods [30] under the same conditions, as the following proposition states.
Corollary 2.
Under the setting of Theorem 2, setting $\alpha=1$ and $b=\mathcal{O}(\sqrt{n})$, the complexity in expectation of Algorithm 1 is $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$.
5 Conclusions
We propose the active acceleration-variance reduction trade-off in single-loop VR methods solving the composite general convex optimization problem. At the core of this idea is the proposed $p_{t}-\alpha_{t}$ correspondence, named the Harmonia technique, which essentially quantitatively determines how much variance reduction is needed to perform certain level of acceleration, or vise versa. The key parameter $\alpha\in[0,1]$ controls the acceleration-variance reduction trade-off and through which essentially controls the trade-off between convergence rate and per-iteration cost continuously in the proposed unifying algorithmic framework.
The continuous nature and flexibility of $\alpha\in[0,1]$ brings several desirable merits. First, several known complexity of some quite different methods are re-discovered under the proposed unifying framework Katyusha-H. Second, under an extra mild condition, Katyusha-H can achieve the near-optimal complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ for $\alpha$ belonging to certain explicitly given interval with the smallest possible mini-batch size $b=1$ and decreasing per-iteration cost in expectation (when $\alpha<1$). Third, setting $\alpha=1$ and $b=\mathcal{O}(\sqrt{n})$, Katyusha-H achieves the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ under standard assumptions without knowledge of the target accuracy, which significantly improves upon previous best $\mathcal{O}(n+n/\sqrt{\epsilon})$ of single-loop VR methods and matches the theoretical lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ up to logarithmic factors.
We believe that the proposed idea and technique may be of general interest beyond the considered problem in this work.
Proof of Lemma 1. We begin by proving the following key inequality:
$\xi\left(\alpha_{t+1}^{2}-\alpha_{t}^{2}\right)\leq\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t},\ \forall t\in\mathbb{N}^{+}.$ (6)
When $\alpha=0$, all $\alpha_{t}$ are equal and positive, hence (6) obviously holds. The cases where $\alpha\neq 0$ are considered as follows. Consider the case where $t\geq 18$, so that $\alpha_{t-1},\alpha_{t},\alpha_{t+1}$ take the form $a_{\alpha}(t-1)^{\alpha},a_{\alpha}t^{\alpha}$, and $a_{\alpha}(t+1)^{\alpha}$, respectively. In this case, it suffices to prove
$\alpha_{t+1}^{2}\leq\alpha_{t}+\alpha_{t-1}^{2},\ \forall t\geq 18,$ (7)
since by (7), we have
$\xi\left(\alpha_{t+1}^{2}-\alpha_{t}^{2}\right)\leq\alpha_{t+1}^{2}-\alpha_{t}^{2}\leq\alpha_{t}-\alpha_{t}^{2}+\alpha_{t-1}^{2},$
where the first inequality follows from $\alpha_{t+1}\geq\alpha_{t}$ and the setting $\xi=1/(bc)$ where $b\in[n],c\geq 2$. Define
$g(t)=a_{\alpha}t^{\alpha}+a_{\alpha}^{2}(t-1)^{2\alpha}-a_{\alpha}^{2}(t+1)^{2\alpha},$
then inequality (7) is equivalent to $g(t)\geq 0$ for $t\geq 18$, which will be proved by a monotonicity argument. The simple case $\alpha=1$ will be discussed separately. We prove $g(17)\geq 0$ for $\alpha\in\left(0,1\right)$ as follows. In the $\alpha\in\left(0,1/2\right]$ case, note that
$g(17)$ $=a_{\alpha}\left[17^{\alpha}+a_{\alpha}\left(16^{2\alpha}-18^{2\alpha}\right)\right]$
$=a_{\alpha}\left[17^{\alpha}+a_{\alpha}\left(16^{\alpha}-18^{\alpha}\right)\left(16^{\alpha}+18^{\alpha}\right)\right]$
$\geq a_{\alpha}\left(16^{\alpha}+18^{\alpha}\right)\left[a_{\alpha}\left(16^{\alpha}-18^{\alpha}\right)+\frac{1}{2}\right]$
$\geq\frac{4+\sqrt{2}}{4}(16^{\alpha}+18^{\alpha})\left[\frac{4+\sqrt{2}}{4}\left(\sqrt{16}-\sqrt{18}\right)+\frac{1}{2}\right],$
where in the first inequality the concavity inequality $17^{\alpha}\geq(1/2)\left(16^{\alpha}+18^{\alpha}\right)$ is used, and the second inequality is by the fact that $r_{1}^{\alpha}-r_{2}^{\alpha}$ is decreasing with respect to (w.r.t.) $\alpha$, where $1<r_{1}<r_{2}$ (this can be shown by taking derivative w.r.t. $\alpha$). Simple calculation shows that $g(17)\geq 0$ in this case. Similarly, in the $\alpha\in\left(1/2,3/4\right]$ case,
$g(17)\geq\frac{1}{3}\left(16^{\alpha}+18^{\alpha}\right)\left[\frac{1}{3}\left(16^{3/4}-18^{3/4}\right)+\frac{1}{2}\right]\geq 0.$
In the $\alpha\in\left(3/4,1\right)$ case,
$g(17)\geq\frac{1}{4}\left(\frac{17}{16}\right)^{\alpha-1}\left(16^{\alpha}+18^{\alpha}\right)\left[\frac{4}{17}\left(\frac{17}{16}\right)^{\alpha}\left(16^{\alpha}-18^{\alpha}\right)+\frac{1}{2}\right].$
Note that the quantity $17^{\alpha}-\left(308/16\right)^{\alpha}$ has the infimum 17/8-17/8 on the interval $\left(3/4,1\right)$. This shows that $g(17)\geq 0$. Define
$g_{1}(t)=t^{\alpha-1}+2a_{\alpha}(t-1)^{2\alpha-1}-2a_{\alpha}(t+1)^{2\alpha-1},$
then $g^{\prime}(t)=a_{\alpha}\alpha g_{1}(t)$ and it suffices to prove $g_{1}(t)\geq 0$ for $t\geq 17$ to complete the monotonicity argument.
Consider the case where $\alpha\in(0,1)$. In this case,
$g_{1}(t)=t^{\alpha-1}\left(1+2a_{\alpha}(t-1)^{\alpha}(1-1/t)^{\alpha-1}-2a_{\alpha}(t+1)^{\alpha}(1+1/t)^{\alpha-1}\right).$
To prove $g_{1}(t)\geq 0$, it suffices to prove
$\frac{1}{2a_{\alpha}}\geq(t+1)^{\alpha}\left(1+1/t\right)^{\alpha-1}-(t-1)^{\alpha}(1-1/t)^{\alpha-1}.$
Note that $(1+1/t)^{\alpha-1}\leq(1-1/t)^{\alpha-1}$, it suffices to prove
$1/(2a_{\alpha})\geq(1-1/t)^{\alpha-1}[(t+1)^{\alpha}-(t-1)^{\alpha}].$
Since $(1-1/t)^{\alpha-1}$ is decreasing w.r.t. $t$, hence it is upper bounded by $(16/17)^{\alpha-1}$ for $t\geq 17$. Define $q_{t}(\alpha)=(t+1)^{\alpha}-(t-1)^{\alpha}$ where $t$ is regarded as a parameter. When $t\geq 17$, taking derivative w.r.t. $\alpha$ gives
$q_{t}^{\prime}(\alpha)=(t+1)^{\alpha}\log{(t+1)}-(t-1)^{\alpha}\log{(t-1)}.$
Since $(t+1)^{\alpha}\geq(t-1)^{\alpha}\geq 0$ and $\log{(t+1)}\geq\log{(t-1)}\geq 0$, we have $q_{t}^{\prime}(\alpha)\geq 0$ for any fixed $t\geq 17$.
Consider the sub-case where $\alpha\in\left(0,1/2\right]$. Fix any $t\geq 17$, we have
$q_{t}(\alpha)\leq\lim_{\alpha\rightarrow(1/2)^{-}}q_{t}(\alpha)=\sqrt{t+1}-\sqrt{t-1}\leq\frac{1}{\sqrt{t-1}}\leq\frac{1}{4}.$
Hence
$\left(1-1/t\right)^{\alpha-1}\left[(t+1)^{\alpha}-(t-1)^{\alpha}\right]\leq(16/17)^{\alpha-1}\cdot(1/4)\leq 1/(2a_{\alpha}),$
where the second inequality is by our choice of $a_{\alpha}=(4+\sqrt{2})/4$, and the inequality desired in this sub-case is proved.
Consider the sub-case where $\alpha\in\left(1/2,3/4\right]$. Fix any $t\geq 17$, we have
$q_{t}(\alpha)\leq\lim_{\alpha\rightarrow(3/4)^{-}}q_{t}(\alpha)=(t+1)^{3/4}-(t-1)^{3/4}\leq\frac{3}{4}.$
Hence
$\left(1-1/t\right)^{\alpha-1}\left[(t+1)^{\alpha}-(t-1)^{\alpha}\right]\leq(16/17)^{\alpha-1}\cdot(3/4)\leq 1/(2a_{\alpha}),$
where the second inequality is by our choice of $a_{\alpha}=1/3$.
Consider the sub-case where $\alpha\in\left(3/4,1\right)$. Fix any $t\geq 17$, we have $q_{t}(\alpha)\leq\lim_{\alpha\rightarrow 1^{-}}q_{t}(\alpha)=2$, so $q_{t}(\alpha)\leq 2$ for $t\geq 17$. Hence
$(1-1/t)^{\alpha-1}\left[(t+1)^{\alpha}-(t-1)^{\alpha}\right]\leq(16/17)^{\alpha-1}\cdot 2\leq 1/(2a_{\alpha}),$
where the second inequality is by our choice of $a_{\alpha}=(1/4)\cdot(17/16)^{\alpha-1}$.
Consider the case where $\alpha=1$. In this case, $g_{1}(t)=1-4a_{\alpha}\geq 0$ since $a_{\alpha}|_{\alpha=1}=1/4$.
Altogether, by monotonicity, we have $g(t)\geq 0$ for $t\geq 18,\alpha\in\left(0,1\right]$. Thus (7) is proved for all $\alpha\in\left(0,1\right]$, which implies that (6) holds for all $t\geq 18$, $\alpha\in\left(0,1\right]$. We then prove (6) for $1\leq t\leq 17$ under all $\alpha\in\left(0,1\right]$ settings (except for $\alpha=0$). When $t=17$, since $\alpha_{18}^{2}\geq\alpha_{17}^{2}$ for $\alpha\in\left(0,1\right]$, by the reasoning at the beginning of this proof, it suffices to prove $\alpha_{18}^{2}\leq\alpha_{17}+\alpha_{16}^{2}$. In the $\alpha\in\left(0,1/2\right]$ sub-case, $\alpha_{18}^{2}=(1+\sqrt{2}/4)^{2}\cdot 18^{2\alpha}\leq 36=\alpha_{16}^{2}$, in the $\alpha\in\left(1/2,3/4\right]$ sub-case, $\alpha_{18}^{2}=(1/9)\cdot 18^{2\alpha}\leq 36$, and in the $\alpha\in\left(3/4,1\right]$ sub-case, $\alpha_{18}^{2}=\left(1/16\right)\cdot\left(17/16\right)^{2\alpha-2}18^{2\alpha}\leq 36=\alpha_{16}^{2}$, this concludes the $t=17$ case for all $\alpha\in\left(0,1\right]$. When $t=16$, $\alpha_{17}^{2}-\alpha_{16}^{2}\leq\alpha_{18}^{2}-\alpha_{16}^{2}\leq 0$, while the RHS of (6) equals $\alpha_{16}\geq 0$, hence the $t=16$ case is proved. Note that (6) obviously holds for $1\leq t\leq 15$, since the left hand side (LHS) equals 0 and the RHS is positive.
Altogether, the key inequality (6) is proved for all $\alpha\in[0,1]$, and its corollaries are listed as follows. First, it follows from (6) that the following inequality holds:
$\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t}\geq 0,\ \forall t\in\mathbb{N}^{+}.$ (8)
Since when $t\in\mathbb{N}^{+}\setminus\{16\}$, we have $\alpha_{t+1}^{2}\geq\alpha_{t}^{2}$, by which one concludes $\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t}\geq 0$, where (6) is used. And when $t=16$, $\alpha_{15}^{2}-\alpha_{16}^{2}+\alpha_{16}=\alpha_{16}\geq 0$. Second, it follows from (6) that the following inequality holds:
$\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t-1}^{2}+\sum_{j=1}^{t-1}\alpha_{j}\geq\xi\alpha_{t}^{2}\geq 0,\ \forall t\in\mathbb{N}^{+}.$ (9)
We show this by mathematical induction. When $t=1$, $\tilde{\alpha}_{0}\geq\xi\alpha_{1}^{2}$ by the setting of $\tilde{\alpha}_{0}$. Suppose (9) holds for some $t\in\mathbb{N}^{+}$. Then
$\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}-\xi\alpha_{t+1}^{2}$ $=\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t-1}^{2}+\sum_{j=1}^{t-1}\alpha_{j}+\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t}-\xi\alpha_{t+1}^{2}$
$\geq\xi\left(\alpha_{t}^{2}-\alpha_{t+1}^{2}\right)+\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t}\geq 0,$
where in the first inequality we use the induction hypothesis and in the second inequality (6) is used. By mathematical induction, (9) is proved. Recalling the definition of the $\{p_{t}\}$ sequence in Algorithm 1, $p_{t}$ can be reformulated as follows
$p_{t}=\frac{\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t}+\xi\alpha_{t}^{2}}{\alpha_{t-1}^{2}-\alpha_{t}^{2}+\alpha_{t}+\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t-1}^{2}+\sum_{j=1}^{t-1}\alpha_{j}}.$ (10)
By (8), (9) and (10), we have
$0\leq p_{t}\leq 1,\ \forall t\in\mathbb{N}^{+}.$ (11)
One easily checks that $\alpha_{17}>1$ for $\alpha=0,\alpha\in\left(0,1/2\right],\alpha\in\left(1/2,3/4\right]$, and $\alpha\in\left(3/4,1\right]$, respectively. In fact, $\alpha_{17}$ allows a uniform deterministic constant lower bound for all $\alpha\in[0,1]$. Specifically, $\alpha_{17}=6$ for $\alpha=0$, $\alpha_{17}\geq 1+\sqrt{2}/4$ for $\alpha\in\left(0,1/2\right]$, $\alpha_{17}\geq 4/3$ for $\alpha\in\left(1/2,3/4\right]$, and $\alpha_{17}\geq 32/17$ for $\alpha\in\left(3/4,1\right]$. By the monotonicity of the $\alpha_{t}$ sequence w.r.t. $t$, and noting that $\alpha_{t}=6$ for $1\leq t\leq 16$, we have $\alpha_{t}>1$ for $\alpha\in[0,1]$, which shows that $\tau_{t}=1/\alpha_{t}\in(0,1)$. Note that $\tau_{t}\leq\max\left\{1/6,\alpha_{17}^{-1}\right\}<1$, by choosing $c=\max\left\{2,b^{-1}\max\left\{6/5,\left(1-\alpha_{17}^{-1}\right)^{-1}\right\}\right\}+1$, we have $1-\tau_{t}-\xi\in(0,1)$. The fact that $\alpha_{17}$ is uniformly bounded away from $1$ guarantees that the constant in $c$ does not explode for any choice of $\alpha\in[0,1]$. Since $b\geq 1$ and $\alpha_{17}\geq 4/3$ for all $\alpha\in[0,1]$, there exists a constant $C>0$ that does not depend on any problem-dependent constant (e.g., $n$) such that [^17] For example, one can take $C=5$.
$c\leq C,\ \forall\alpha\in[0,1].$ (12)
And by the choice of $c$, we also have $\xi\in(0,1)$. Altogether, the following inequalities are ensured
$\tau_{t}\in(0,1),\ \xi\in(0,1),\ (1-\xi-\tau_{t})\in(0,1),\ \forall t\in\mathbb{N}^{+}.$ (13)
[…]
The following lemma mimics Lemma $2.4$ of [8].
Lemma 2.
In Algorithm 1, the variance of the mini-batch version of the SVRG estimator $\tilde{g}_{t+1}$ can be bounded as follows:
$\mathbb{E}_{J_{t}}\left[\|\tilde{g}_{t+1}-\nabla f(x_{t+1})\|^{2}\right]\leq\frac{2L}{b}\left(f(w_{t})-f(x_{t+1})-\langle\nabla f(x_{t+1}),w_{t}-x_{t+1}\rangle\right).$
Proof.
For a given iteration $t\in\mathbb{N}^{+}$, define $X_{j}=\nabla f_{j}(x_{t+1})-\nabla f_{j}(w_{t})$. Then
$\mathbb{E}_{J_{t}}\sum_{j\in J_{t}}X_{j}=\sum_{\text{all possible}\ J_{t}}\frac{1}{C_{n}^{b}}\sum_{j\in J_{t}}X_{j}=\frac{1}{C_{n}^{b}}\sum_{j=1}^{n}C_{n-1}^{b-1}X_{j}=\frac{b}{n}\sum_{j=1}^{n}X_{j},$ (14)
where $C_{n}^{m}=n!/\left[m!(n-m)!\right]$. Hence
$\mathbb{E}_{J_{t}}\|\tilde{g}_{t+1}-\nabla f(x_{t+1})\|^{2}$ $=\mathbb{E}_{J_{t}}\|\frac{1}{b}\sum_{j\in J_{t}}X_{j}+\nabla f(w_{t})-\nabla f(x_{t+1})\|^{2}$
$=\mathbb{E}_{J_{t}}\|\frac{1}{b}\sum_{j\in J_{t}}X_{j}-\mathbb{E}_{J_{t}}\frac{1}{b}\sum_{j\in J_{t}}X_{j}\|^{2}$
$\leq\frac{1}{b^{2}}\mathbb{E}_{J_{t}}\|\sum_{j\in J_{t}}X_{j}\|^{2}$
$=\frac{1}{bn}\sum_{j=1}^{n}\|\nabla f_{j}(x_{t+1})-\nabla f_{j}(w_{t})\|^{2}$
$\leq\frac{2L}{b}\left(f(w_{t})-f(x_{t+1})-\langle\nabla f(x_{t+1}),w_{t}-x_{t+1}\rangle\right),$
where in the second equality we applied (14), the third equality can be derived in a way similarly to (14) and we refer to Lemma $18$ and Lemma $19$ of [30], in the last inequality the following inequality is applied:
$\frac{1}{2L}\|\nabla g(x)-\nabla g(y)\|^{2}\leq g(x)-g(y)-\langle\nabla g(y),x-y\rangle,\ \forall x,y\in\mathbb{R}^{d},$
where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a L-smooth convex function, and this inequality is a standard tool which can be found in, e.g., [43]. ∎
With Lemma 1 and Lemma 2, the convergence theorem can be proved as follows.
Proof of Theorem 1. Note that
$\alpha_{t}\left(f(x_{t+1})-f(x^{*})\right)$ $\leq\alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-x^{*}\rangle$
$=\alpha_{t}\langle\nabla f(x_{t+1}),z_{t}-x^{*}\rangle+\alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-z_{t}\rangle$
$=\mathbb{E}_{J_{t}}\alpha_{t}\langle\tilde{g}_{t+1},z_{t}-x^{*}\rangle+\alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-z_{t}\rangle$
$=\mathbb{E}_{J_{t}}\alpha_{t}\langle\tilde{g}_{t+1},z_{t}-z_{t+1}\rangle+\mathbb{E}_{J_{t}}\alpha_{t}\langle\tilde{g}_{t+1},z_{t+1}-x^{*}\rangle$
$+\alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-z_{t}\rangle.$ (15)
where in the first inequality we applied the convexity of $f$, and the second equality is by the unbiasedness of the estimator $\tilde{g}_{t+1}$. The three terms in the RHS of (15) are separately bounded as follows. For the first term,
$\alpha_{t}\langle\tilde{g}_{t+1},z_{t}-z_{t+1}\rangle$ $=\alpha_{t}\langle\nabla f(x_{t+1}),z_{t}-z_{t+1}\rangle+\alpha_{t}\langle\tilde{g}_{t+1}-\nabla f(x_{t+1}),z_{t}-z_{t+1}\rangle$
$=\alpha_{t}^{2}\langle\nabla f(x_{t+1}),x_{t+1}-y_{t+1}\rangle+\alpha_{t}^{2}\langle\tilde{g}_{t+1}-\nabla f(x_{t+1}),x_{t+1}-y_{t+1}\rangle$
$\leq\alpha_{t}^{2}\left[f(x_{t+1})-f(y_{t+1})+\frac{L}{2}\|x_{t+1}-y_{t+1}\|^{2}\right]+\frac{\alpha_{t}^{2}}{2Lc}\|\tilde{g}_{t+1}-\nabla f(x_{t+1})\|^{2}$
$+\frac{\alpha_{t}^{2}Lc}{2}\|x_{t+1}-y_{t+1}\|^{2}$
$\leq\alpha_{t}^{2}\left[f(x_{t+1})-f(y_{t+1})+\frac{L}{2}\|x_{t+1}-y_{t+1}\|^{2}\right]+\frac{\alpha_{t}^{2}Lc}{2}\|x_{t+1}-y_{t+1}\|^{2}$
$+\frac{\alpha_{t}^{2}}{bc}\left[f(w_{t})-f(x_{t+1})-\langle\nabla f(x_{t+1}),w_{t}-x_{t+1}\rangle\right],$ (16)
where in the second equality the relation $z_{t}-z_{t+1}=\alpha_{t}(x_{t+1}-y_{t+1})$ is used, in the first inequality the $L$-smoothness of $f$ and the Young inequality are used, and the last inequality is by Lemma 2. For the term $\alpha_{t}\langle\tilde{g}_{t+1},z_{t+1}-x^{*}\rangle$, we have
$\alpha_{t}\langle\tilde{g}_{t+1},z_{t+1}-x^{*}\rangle$ $=\langle\frac{1}{\eta}(z_{t}-z_{t+1})-\alpha_{t}\bar{g}_{t+1},z_{t+1}-x^{*}\rangle$
$=\frac{1}{\eta}\langle z_{t}-z_{t+1},z_{t+1}-x^{*}\rangle-\alpha_{t}\langle\bar{g}_{t+1},z_{t+1}-x^{*}\rangle$
$\leq\frac{1}{2\eta}\left(\|z_{t}-x^{*}\|^{2}-\|z_{t+1}-x^{*}\|^{2}-\|z_{t}-z_{t+1}\|^{2}\right)$
$+\alpha_{t}\left(l(x^{*})-l(z_{t+1})\right).$ (17)
where $\bar{g}_{t+1}$ is a subgradient of $l$ at $z_{t+1}$, and in the first equality we use the equivalent update formula [^18] This is derived from the optimality of $z_{t+1}$ in the optimization problem $z_{t+1}=\arg\min_{z\in\mathbb{R}^{d}}\left\{\frac{1}{2\alpha_{t}\eta}\|z-z_{t}\|^{2}+\langle\tilde{g}_{t+1},z\rangle+l(z)\right\},$ which is a standard result and can be found in, e.g., [44]. of $z_{t+1}$, i.e., $z_{t+1}=z_{t}-\alpha_{t}\eta\left(\tilde{g}_{t+1}+\bar{g}_{t+1}\right)$, in the last inequality we use the convexity of $l$ and the three-point property $\langle a-b,b-c\rangle=(1/2)\cdot(\|a-c\|^{2}-\|a-b\|^{2}-\|b-c\|^{2}),\ \forall a,b,c\in\mathbb{R}^{d}$. Recall $x_{t+1}=\tau_{t}z_{t}+\xi w_{t}+(1-\xi-\tau_{t})y_{t}$ in Algorithm 1 and the setting $\tau_{t}\alpha_{t}=1$, we have
$x_{t+1}-z_{t}=\alpha_{t}\xi(w_{t}-x_{t+1})+\alpha_{t}(1-\xi-\tau_{t})(y_{t}-x_{t+1}).$ (18)
Then for the term $\alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-z_{t}\rangle$, we have
$\quad\ \alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-z_{t}\rangle$
$=\alpha_{t}^{2}\xi\langle\nabla f(x_{t+1}),w_{t}-x_{t+1}\rangle+\alpha_{t}^{2}(1-\xi-\tau_{t})\langle\nabla f(x_{t+1}),y_{t}-x_{t+1}\rangle$
$\leq\alpha_{t}^{2}\xi\langle\nabla f(x_{t+1}),w_{t}-x_{t+1}\rangle+\alpha_{t}^{2}(1-\xi-\tau_{t})(f(y_{t})-f(x_{t+1})),$ (19)
where the first equality is by (18), the last inequality is by the convexity of $f$ and Lemma 1 which ensures $1-\xi-\tau_{t}\geq 0$. Taking expectation and combining (15), (16), (17), and (19), we have
$\quad\ \alpha_{t}\mathbb{E}\left(f(x_{t+1})-f(x^{*}))\right)$
$\leq\alpha_{t}^{2}\mathbb{E}\left[f(x_{t+1})-f(y_{t+1})+\frac{L}{2}\|x_{t+1}-y_{t+1}\|^{2}\right]+\frac{\alpha_{t}^{2}Lc}{2}\mathbb{E}\|x_{t+1}-y_{t+1}\|^{2}$
$+\alpha_{t}^{2}\xi\mathbb{E}\left[f(w_{t})-f(x_{t+1})-\langle\nabla f(x_{t+1}),w_{t}-x_{t+1}\rangle\right]$
$+\frac{1}{2\eta}\mathbb{E}\left(\|z_{t}-x^{*}\|^{2}-\|z_{t+1}-x^{*}\|^{2}-\|z_{t}-z_{t+1}\|^{2}\right)+\alpha_{t}\mathbb{E}\left(l(x^{*})-l(z_{t+1})\right)$
$+\alpha_{t}^{2}\xi\mathbb{E}\langle\nabla f(x_{t+1}),w_{t}-x_{t+1}\rangle+\alpha_{t}^{2}(1-\xi-\tau_{t})\mathbb{E}\left(f(y_{t})-f(x_{t+1})\right).$ (20)
Note again that $z_{t+1}-z_{t}=\alpha_{t}(y_{t+1}-x_{t+1})$ by the $y_{t+1}$ update in Algorithm 1, inequality (20) becomes
$\alpha_{t}^{2}\mathbb{E}\left(f(y_{t+1})-F(x^{*})\right)$ $\leq\alpha_{t}^{2}(1-\xi-\tau_{t})\mathbb{E}\left(f(y_{t})-F(x^{*})\right)+\alpha_{t}^{2}\xi\mathbb{E}\left(f(w_{t})-F(x^{*})\right)$
$-\alpha_{t}\mathbb{E}l(z_{t+1})+\alpha_{t}^{2}\left(\frac{L}{2}+\frac{Lc}{2}-\frac{1}{2\eta}\right)\mathbb{E}\|x_{t+1}-y_{t+1}\|^{2}$
$+\frac{1}{2\eta}\mathbb{E}\|z_{t}-x^{*}\|^{2}-\frac{1}{2\eta}\mathbb{E}\|z_{t+1}-x^{*}\|^{2}.$ (21)
By update formulae of $x_{t+1}$ and $y_{t+1}$ in Algorithm 1, we have
$y_{t+1}=\tau_{t}z_{t+1}+\xi w_{t}+(1-\xi-\tau_{t})y_{t},$
it follows (by Lemma 1, $\tau_{t}$, $\xi$, and $(1-\xi-\tau_{t})$ all belong to $(0,1)$)
$l(y_{t+1})\leq\tau_{t}l(z_{t+1})+\xi l(w_{t})+(1-\xi-\tau_{t})l(y_{t}).$ (22)
Note that $L/2+Lc/2-1/(2\eta)\leq 0$ since $\eta\leq 1/(cL+L)$. Then combining (21) and (22), we arrive at
$\alpha_{t}^{2}\mathbb{E}\left(F(y_{t+1})-F(x^{*})\right)$ $\leq\alpha_{t}^{2}(1-\xi-\tau_{t})\mathbb{E}\left(F(y_{t})-F(x^{*})\right)+\alpha_{t}^{2}\xi\mathbb{E}\left(F(w_{t})-F(x^{*})\right)$
$+\frac{1}{2\eta}\mathbb{E}\|z_{t}-x^{*}\|^{2}-\frac{1}{2\eta}\mathbb{E}\|z_{t+1}-x^{*}\|^{2}.$ (23)
By the tower property of conditional expectation and the probabilistic update formula of $w_{t+1}$ in Algorithm 1, we have
$\mathbb{E}\left(F(w_{t+1})-F(x^{*})\right)\leq(1-p_{t})\mathbb{E}\left(F(w_{t})-F(x^{*})\right)+p_{t}\mathbb{E}\left(F(y_{t})-F(x^{*})\right).$ (24)
For $t\in\mathbb{N}$, define
$\mathcal{L}_{t+1}=$ $\alpha_{t}^{2}\mathbb{E}\left(F(y_{t+1})-F(x^{*})\right)+\left(\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\right)\mathbb{E}\left(F(w_{t+1})-F(x^{*})\right)$
$+\frac{1}{2\eta}\mathbb{E}\|z_{t+1}-x^{*}\|^{2}.$
Multiplying (24) by $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}$ and combining the resultant inequality with (23), we obtain (note that $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq 0$ by Lemma 1)
$\mathcal{L}_{t+1}\leq\mathcal{L}_{t},\ \forall t\in\mathbb{N},$ (25)
[…]
Proof of Theorem 2. We divide the proof into cases where the $\{\alpha_{t}\}$ sequence has different growing rates, which is determined by $\alpha$ in Lemma 1. In all cases, we always investigate the convergence of the quantity $\mathbb{E}\left(F(w_{t})-F(x^{*})\right)$.
In the case where $\alpha=0$, we have $\alpha_{t}\equiv 6$ for all $t\in\mathbb{N}$. In this case, we have
$p_{t}=\frac{36\xi+6}{\tilde{\alpha}_{0}+6t}.$
By (2), Algorithm 1 needs $T=\mathcal{O}(1/\epsilon)$ iterations to reach an $\epsilon$-accurate solution of (1). In expectation, Algorithm 1 calls IFO $\mathcal{O}(b+np_{t})$ times per iteration. Hence the IFO complexity in expectation $\mathbb{E}C_{\text{comp}}$ is $\mathcal{O}(b/\epsilon+n\sum_{t=1}^{\lceil 1/\epsilon\rceil}p_{t})$, where the dominant term can be further bounded as follows
$\frac{b}{\epsilon}+n\sum_{t=1}^{\lceil 1/\epsilon\rceil}p_{t}$ $\leq\frac{b}{\epsilon}+\frac{n\left(36\xi+6\right)}{6}\sum_{t=1}^{\lceil 1/\epsilon\rceil}\frac{1}{t}$
$=\frac{b}{\epsilon}+n\left(6\xi+1\right)+n\left(6\xi+1\right)\sum_{t=2}^{\lceil 1/\epsilon\rceil}\frac{1}{t}$
$\leq\frac{b}{\epsilon}+7n+7n\log\left(\lceil 1/\epsilon\rceil\right),$
where the last inequality is by $\xi\in(0,1)$ and $\sum_{t=2}^{\lceil 1/\epsilon\rceil}(1/t)\leq\int_{1}^{\lceil 1/\epsilon\rceil}(1/t)\text{d}t$. Hence in this case, the optimal choice of the mini-batch size is $b^{*}=1$, and the IFO complexity in expectation is $\mathcal{O}\left(n\log(1/\epsilon)+1/\epsilon\right)$.
In the case where $\alpha\in(0,1]$, we have
$\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}$ $=\tilde{\alpha}_{0}+132-a_{\alpha}^{2}t^{2\alpha}+a_{\alpha}\sum_{j=17}^{t}j^{\alpha}$
$\geq\tilde{\alpha}_{0}+132-\frac{a_{\alpha}}{\alpha+1}16^{\alpha+1}-a_{\alpha}^{2}t^{2\alpha}+\frac{a_{\alpha}}{\alpha+1}t^{\alpha+1},$
where the inequality is by $\sum_{j=17}^{t}j^{\alpha}\geq\int_{16}^{t}u^{\alpha}\text{d}u$. Note that $132\geq\left(a_{\alpha}16^{\alpha+1}\right)/(\alpha+1)$ for all $\alpha\in\left(0,1\right]$. For $\alpha\in(0,1)$, define
$T_{\alpha}=\left[2a_{\alpha}(\alpha+1)\right]^{\frac{1}{1-\alpha}},$
then we have $a_{\alpha}^{2}t^{2\alpha}\leq\left(a_{\alpha}t^{\alpha+1}\right)/(2\alpha+2)$ for all $t\in\mathbb{N}^{+}\cap\left[T_{\alpha},+\infty\right)$. One easily checks that $T_{\alpha}\leq 17$ for all $\alpha\in(0,1)$. Moreover, when $\alpha=1$, we have $a_{\alpha}|_{\alpha=1}=1/4$, and $a_{\alpha}^{2}t^{2\alpha}+\left(a_{\alpha}t^{\alpha+1}\right)/(\alpha+1)=\frac{t^{2}}{16}$. For $\alpha\in\left(0,1\right]$, define
$\tilde{a}_{\alpha}=\begin{cases}\frac{a_{\alpha}}{2\alpha+2}&\text{when}\ \alpha\in(0,1)\\ \frac{1}{16}&\text{when}\ \alpha=1\end{cases},$
then for any $\alpha\in\left(0,1\right]$, when $t\geq 17$, we have $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq\tilde{a}_{\alpha}t^{\alpha+1}$. Since $T_{\alpha}$ admits a uniform constant upper bound $17$ for any $\alpha\in\left(0,1\right]$, one can find a $\tilde{\tilde{a}}_{\alpha}>0$ (which is free from problem dependent parameters and possibly smaller than $\tilde{a}_{\alpha}$) so that $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq\tilde{\tilde{a}}_{\alpha}t^{\alpha+1}$ for all $t\in\mathbb{N}^{+}$. Hence by Theorem 1, Algorithm 1 needs $T=\mathcal{O}\left(1/\epsilon^{1/\left(\alpha+1\right)}\right)$ iterations to reach an $\epsilon$-accurate solution of (1). To estimate $\mathbb{E}C_{\text{comp}}=\left(b/\epsilon^{1/\left(\alpha+1\right)}+n\sum_{t=1}^{\lceil 1/\epsilon^{1/\left(\alpha+1\right)}\rceil}p_{t}\right)$, the update probability $p_{t}$ is bounded as follows
$p_{t}$ $\leq\tilde{a}_{\alpha}^{-1}\frac{a_{\alpha}^{2}(t-1)^{2\alpha}-(1-\xi)a_{\alpha}^{2}t^{2\alpha}+a_{\alpha}t^{\alpha}}{t^{\alpha+1}}$
$=\tilde{a}_{\alpha}^{-1}\frac{a_{\alpha}^{2}\left[(t-1)^{2\alpha}-t^{2\alpha}\right]+a_{\alpha}^{2}\xi t^{2\alpha}+a_{\alpha}t^{\alpha}}{t^{\alpha+1}},$ (26)
where $\xi=1/(bc)$ by our setting and $t\geq 17$. Note that although for $t<17$ the corresponding $p_{t}$ is not upper bounded carefully, we can always upper bound it by $1$, which at most leads to an addend being constant multiple of $n$ in the complexity, i,e., $\mathcal{O}(n)$. For $t\geq 17$, the inequality (26) implies
$p_{t}\leq\tilde{a}_{\alpha}^{-1}a_{\alpha}^{2}\xi\frac{1}{t^{1-\alpha}}+\tilde{a}_{\alpha}^{-1}a_{\alpha}\frac{1}{t}.$
Note that
$Q_{1}(\epsilon,\alpha)=\sum_{t=17}^{\lceil 1/\epsilon^{1/\left(\alpha+1\right)}\rceil}\frac{1}{t^{1-\alpha}}$ $\leq\int_{16}^{\lceil 1/\epsilon^{1/\left(\alpha+1\right)}\rceil}\frac{1}{t^{1-\alpha}}\text{d}t=\mathcal{O}\left(\frac{1}{\alpha}\frac{1}{\epsilon^{\frac{\alpha}{\alpha+1}}}\right),$ (27)
$Q_{2}(\epsilon,\alpha)$ $=\sum_{t=17}^{\lceil 1/\epsilon^{1/\left(\alpha+1\right)}\rceil}\frac{1}{t}=\mathcal{O}\left(\frac{\log(1/\epsilon)}{\alpha+1}\right).$ (28)
The quantity $Q_{1}(\epsilon,\alpha)$ is essentially upper bounded by $(1/\alpha)\cdot(1/\epsilon^{(\alpha/(\alpha+1))})$, which explodes when $\alpha\rightarrow 0^{+}$, hence the $\alpha\in\left(0,1/2\right]$ case is further handled as follows. Without loss of generality, assume $\epsilon\in(0,1)$, then for $\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right]$,
$Q_{1}(\epsilon,\alpha)=\sum_{t=17}^{\lceil 1/\epsilon^{1/(\alpha+1)}\rceil}\frac{t^{\alpha}}{t}\leq\sum_{t=17}^{\lceil 1/\epsilon\rceil}\frac{\left(\lceil 1/\epsilon\rceil\right)^{\hat{\alpha}}}{t}=\mathcal{O}\left(\log\left(1/\epsilon\right)\right).$
For $\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1/2\right]$, the quantity $1/\alpha$ in the RHS of (27) can be upper bounded by $\max\left\{1/\hat{\alpha},10\right\}$, which is at the order of $\mathcal{O}(\log(1/\epsilon))$ [^19] For $\alpha\in\left(0,1/2\right]$ which is not very close to 0, e.g., $\alpha=1/4$, one can directly and safely apply (27).. Then the complexity in expectation is
$\mathbb{E}C_{\text{comp}}=\begin{cases}\mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+n\log(1/\epsilon)\right),\quad&\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right],\\ \mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+\frac{n}{b}\frac{1}{\alpha}\frac{1}{\epsilon^{\alpha/(\alpha+1)}}+n\log(1/\epsilon)\right),\quad&\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right].\end{cases}$
[…]